8. Find the solution to the given systems. (20pts) 0 3 (a) X = X X(0) = =) 1 2dx (b) = 5x+y dt dy = -2x+3y dt
Answer #1

8. a) x 12 X -12 ; X(0) = (6) 3 5 e) -1/2 ) Let A= S 12 -1/2 Characteristie polynomial of A, 1 - Ź -> (3--)(-1-7)= 0 => a =b d2 dt de 5x+Y du i =-22+ 3y => 6-10%) = * = ar Characteristic equation of A, 5-2 1 = - 2 3-2 2²8 15-8ata ²+2=0 82 +17 = 0 »4t Gettrost X(t)= Gent cost sint tie -cost-sint cost-sint Hence the general solution is 4tr sint cost-sini the sint YCZ) = Ge